# Enabling Physical Localization of Uncooperative Cellular Devices

### v2

@@ 3.2 Threat Model
We consider an unprivileged attacker, who cannot control the operation of the network infrastructure (e.g., eNB or core networks) or run its own code on target devices. This scenario applies both to authorities (such as police), and attackers. In particular, this setting does also not rely on optional features that are not widely supported in UEs and infrastructure (e.g., 3GPP positioning features).
We assume that the attacker knows an online identity (e.g., phone number) of the victim and the cell information which the victim is camping on. The cell information can be obtained in advance by exploiting the existing presence test [9, 10, 12, 25, 33, 35, 46]. Alternatively, the police agency can request cell information from mobile carriers for their investigation.
@@
Instead of accessing the network and UE directly, the attacker employs multiple devices equipped with directional antennas (i.e., AoA) to measure the direction of the target UE’s uplink signal. Using these devices, the attacker aims to determine the physical location (e.g., 2D-coordinated) of the victim’s UE by measuring the direction of uplink signals at multiple points. In addition, these devices have the capability to passively monitor both uplink and downlink LTE signals. This capability can be achieved using an open-source LTE eavesdropper, such as LTESniffer [24], with off-the-shelf software-defined radio devices (e.g., USRP [51]). One unique capability of the devices is the ability to actively inject malicious LTE signals (i.e., signal overshadowing [19, 54]) on both downlink and uplink channels. Upon obtaining the victim’s uplink scheduling information using these devices, the attacker can employ a directional antenna to measure the direction of the uplink signal source, detecting the strongest signal power. Utilizing Angle of Arrival (AoA) measurements from multiple locations, the attacker performs a multiangulation process to ascertain the precise coordinates of the victim. Alternatively, the attacker can progressively refine the victim’s location by continuously tracking the direction of the uplink signal source (i.e., iterative direction finding).
TABLE I: Uplink transmission power depending on the distance between the eNB and the UE.
Distance (m) 10 20 30 40 50 60 70 80 90 100 110
Tx PWR (dBm) -7.04 -10.64 0.49 4.74 5.05 4.33 7.65 7.45 7.2 7.39 7.56
RSRP (dBm) -62.37 -62.63 -73.43 -76.66 -78.91 -81.69 -85.36 -85.51 -86.45 -84.23 -87.6
@@
Additionally, among various cellular localization techniques, we use AoA measurements of uplink signals as a fundamental technique for physical localization. Other techniques such as TA, ToA, and TDoA have some limitations in employing fine-grained localization of uncooperative UEs in real-world environments. ToA-based cellular localization [31] requires an additional process for fingerprinting the victim’s device to account for hardware errors. This is needed to accurately calculate the arrival signal timing to accurately estimate the victim’s location. In addition, TA-based localization [48] shows only sub-cell-level localization performance with 78m granularity.
3.3 Challenges of Uncooperative Localization
@@
Uncooperative localization attack in real-world environments needs to address the following three challenges:
@@
To confirm the existence of such challenges in operational cellular networks, their impact will be explored using an LTE channel sniffer (Appendix § B.1).
@@
C1: High dependency on victim’s uplink signal. Recall that the primary property exploited by the localization attacker is the monitoring of the uplink signal generated by the victim’s UE. From this perspective, the attacker can initiate localization only when the victim generates uplink traffic. In other words, the localization attack is challenging when the victim is silent (i.e., not transmitting any signal). In such instances, the attacker is compelled to passively wait for the victim’s uplink traffic using their sniffer, as they lack access to either the eNB or the victim UE.
@@
Additionally, it is important to note that RNTI is a temporary identifier that changes frequently (approximately every 15 to 30 seconds [8, 30, 43]). This implies that when a UE is not actively utilizing any services for a specific duration after connection, its radio connection is terminated by the eNB, and the previously assigned RNTI expires. If the UE later requests LTE services from the eNB, a new RNTI value will be assigned for radio communication. The attacker may face a challenge in identifying the victim’s location if their radio connection and assigned RNTI expire during the localization process. This necessitates the attacker to patiently wait for the victim to establish a new radio connection, after which they must track the victim’s new RNTI to resume the localization process.
@@
C2: Undetectable victim’s signal. The successful measurement of arrival signals at the attacker’s sniffer is crucial for executing the localization attack. If the sniffer fails to detect the victim’s uplink signal, estimating the victim’s location becomes challenging. This issue arises because the UE does not consistently transmit its signal with high power to the eNB. Notably, when the UE is situated in close proximity to the eNB, weak uplink transmission power is used. The UE dynamically adjusts its uplink transmission power to compensate for signal path loss, influenced by factors such as distance and obstacles, in coordination with the eNB [2, 37]. For instance, Table I demonstrates how the UE adjusts its transmission power based on both distance and Reference Signals Received Power (RSRP).
@@
The transmission power control between the eNB and UE creates a shadow area where the victim’s signal remains undetectable by the sniffer. To illustrate this shadow area in real-world environments, we examine how the UE’s position influences the sniffer’s signal detection. In our experiment, the attacker’s sniffer, positioned 80m away from the victim, successfully detected uplink signals from victims located 90m away from the eNB. However, when the victim’s UE was only 10m away from the eNB, the sniffer failed to detect uplink signals at the same position (refer to Appendix § B.3 for detailed experimental setup). Unfortunately, due to the attacker’s lack of permission to control networks and UE, they are unable to legitimately instruct the UE to increase its transmission power when encountering the shadow area.
C3: Differentiating the victim from the repeater.
Figure 3: Impact of cellular repeater in localization attack.
@@
A cellular repeater is a device designed to amplify both downlink and uplink cellular signals, reducing signal path loss between the eNB and UE (i.e., improving cellular channel quality). As a consequence of the repeater’s operation minimizing signal path loss, the UE tends to transmit uplink data with lower signal power. Simultaneously, the repeater transmits (i.e., relays) the data with amplified, higher signal power. It is widely deployed to enhance the service coverage of cellular carriers. However, repeaters pose a challenge for attackers in determining the direction of the strongest signal.
Figure 4: Overview of UMA.
@@ 3.4 Approach: UMA
We present an uncooperative multiangulation attack (UMA), which resolves the key challenges of uncooperative localization (§ 3.3). UMA leverages AoA measurements with directional antennas, composed of six steps (Figure 4). These steps include RNTI acquisition using the victim’s online identity (①–②), manipulation of uplink scheduling (③), boosting the target’s uplink signal strength (④), and measurement of the AoA of the uplink signal (⑤–⑥).
@@
In particular, we come up with two approaches to address the challenges of uncooperative localization; the scheduling manipulation attack and the power boosting attack. The scheduling manipulation attack is specifically designed to address challenge C1, allowing the attacker to acquire the unexpired victim’s RNTI and ensure continuous uplink transmission. To overcome challenges C2 and C3, we employ the power boosting attack, enabling the attacker to increase the victim’s uplink transmission power and distinguish the victim’s signal from the repeater’s. Both attacks exploit vulnerable protocols in LTE specification, which lack implemented security protection measures. By utilizing these techniques, UMA effectively addresses the key challenges associated with uncooperative localization.
@@
Overall procedures of UMA. An UMA attacker follows a systematic process illustrated in  Figure 4. The first two steps are to obtain the victim’s RNTI associated with their phone number (§ 4.1). This is accomplished by sending silent Short Message Service (SMS) messages (①) and monitoring traffic patterns on the downlink channel (②). Note that this RNTI is a key identifier to figure out the victim’s uplink scheduling. Next, the attacker proceeds with a scheduling manipulation attack (③); it transmits fake scheduling requests, and reports fake transmission buffer status to the eNB, posing as the victim UE with its RNTI (§ 4.2). This attack ensures the victim’s radio connection remains active, and their RNTI stays unexpired. Through the scheduling manipulation attack, the attacker forces the victim UE to generate uplink traffic even when there is no data available for transmission in its buffer. Subsequently, the attacker conducts a power boosting attack (④) by injecting manipulated commands for transmission power control logic into the victim’s UE, disguising as the eNB. This attack forces the victim UE to increase its uplink transmission power up to a maximum level (23dBm), overcoming practical challenges caused by the undetectable victim’s uplink signal and the presence of a cellular repeater in the radio network (§ 4.3). After these active operations, the attacker obtains the victim’s scheduling information broadcasted over DCI 0 by monitoring the downlink channel (⑤). With this information, the attacker identifies the victim’s uplink signals on the uplink channel. Then, the attacker measures the AoA of the victim’s uplink signal by rotating a directional antenna mounted on the sniffer (⑥). Ultimately, the attacker can achieve physical localization by tracking down the victim’s physical location iteratively or employing multiangulation with multiple sniffers.
4 Design of UMA
@@
In this section, we present the design of our approaches in a sequential manner. Firstly, we introduce a simple yet feasible method, inspired by previous works, for acquiring the victim’s RNTI using only their online identity (§ 4.1). Next, we describe how an unprivileged attacker compels the victim to generate uplink traffic and maintain the victim’s RNTI (§ 4.2). Finally, we illustrate how the attacker compels the victim to increase their uplink transmission power and overcome cellular repeaters (§ 4.3).
4.1 From Online Identity to Radio Identifier
@@
For an attacker armed solely with the target’s online identity (e.g. phone number, WhatsApp ID), the initial step in UMA is to acquire the target’s temporary identifier used in the radio network (RNTI). The eNB broadcasts uplink scheduling information to each UE, designating a recipient UE by using the RNTI. Consequently, an unprivileged attacker seeking to monitor and ascertain the uplink signal of the victim UE must be aware of the victim UE’s RNTI. Since the RNTI is intentionally designed to be decoupled from the online identity, acquiring the RNTI becomes essential to establish the link between the victim’s online identity and its corresponding RNTI.
@@
Inspired by prior works [25, 27, 30, 33, 35, 43, 46], our approach to acquiring the RNTI involves correlating voice/SMS traffic pattern directed toward the victim with monitored downlink packets identified by the recipient’s RNTI. To this end, we generate successive silent voice calls/SMSes. Also, we leverage our observation that traffic related to control messages generated by incoming voice calls and SMS is distinguishable from user-plane data traffic.
@@
RNTI acquisition is conducted through the following three steps. ➀ Given the victim’s online identity, the attacker makes silent voice calls or sends silent SMSes to the victim following a specific traffic pattern (e.g., multiple times with a constant time gap). ➁ The attacker monitors the downlink traffic of SMS and voice calls. Specifically, they monitor the dedicated channels, called bearers, used for sending control messages of voice calls and SMSes. Note that the attacker is able to monitor the targeted traffic pattern clearly as each bearer is designed to deliver a different type of data. For example, the messages for controlling the voice calls are delivered over a bearer named Data Radio Bearer (DRB) 1, and user-plane traffic and SMS control messages are flooded through DRB 2 and Signalling Radio Bearer (SRB), respectively. ➂ The attacker examines radio connections and finds out the connection showing the intended traffic pattern. Finally, the attacker determines the RNTI of that connection as belonging to the victim.
4.2 Manipulating Victim’s Uplink Scheduling
@@
One of the key pillars of UMA lies in its ability to a) compel the victim to generate uplink transmission and b) retain the acquired victim’s radio identifier (RNTI). To achieve this, we introduce a novel approach known as the Scheduling Manipulation Attack. Essentially, the Scheduling Manipulation Attack injects a crafted message into the eNB, prompting the eNB to schedule resources for uplink transmission to the victim UE and triggering the UE to transmit uplink data. This strategic action prevents the UE from entering an idle state, ensuring the continuity of the acquired victim UE’s radio identifier (i.e., RNTI) obtained from the victim’s online identity (§ 4.1).
4.2.1 Uplink Scheduling Procedure
@@
We begin by outlining the control mechanism governing a UE’s uplink data transmission. Essentially, the UE is granted permission to transmit uplink data only when it receives uplink grants allocated by the eNB. The UE follows a defined procedure to secure uplink grants for data transmission to the eNB. Initially, the UE transmits a scheduling request [2], signaling to the eNB its intention to transmit uplink data. Upon receiving the scheduling request, the eNB grants approval for the UE’s Buffer Status Reporting (BSR) [4], a mechanism used to report the size of uplink data. Subsequently, the UE initiates the BSR, leading to the allocation of sufficient uplink grants by the eNB, allowing the UE to transmit its data.
@@
Scheduling Request is a PHY layer message sent from the UE to the eNB over Physical Uplink Control Channel (PUCCH). To transmit a scheduling request to the eNB, the UE uses schedulingRequestConfig [2, 5] previously shared via the RRC connection setup message. schedulingRequestConfig notifies each UE of which resource should be used to transmit scheduling requests.
@@
4.2.2 Design of Scheduling Manipulation Attack
Figure 5: Procedure of scheduling manipulation attack.

### v3

@@ 3.2. Threat Model
We consider a scenario where an attacker aims to determine the physical location of target devices that are uncooperative in the localization process, such as not reporting their location or controlling their uplink transmission. In this situation, the attacker cannot directly control the network infrastructure (e.g., eNB or core networks) or run their own code on target devices. This scenario applies to both authorities (e.g., police) and malicious actors. Importantly, this setting does not rely on optional features that are not widely supported in UEs and infrastructure, ensuring broader applicability.
We assume that the attacker knows an online identity (e.g., phone number) of the victim and the cell information which the victim is camping on. The cell information can be obtained in advance by exploiting the existing presence test (Baek et al., 2023; Cheng et al., 2023; Hong et al., 2018; Kune et al., 2012; Ludant et al., 2023; Shaik et al., 2016; Bitsikas et al., 2023) or requesting cell information from mobile carriers for investigation (Downing, 2016). Law enforcement agencies, in particular, can obtain the target’s network identifier and cell information with carrier assistance. Nevertheless, this assistance is insufficient for reliable fine-grained physical localization of uncooperative devices, as mentioned in § 3.1. Therefore, in this scenario, carriers are considered semi-cooperative with law enforcement’s localization efforts.
@@
The attacker employs multiple devices equipped with directional antennas (i.e., AoA) to measure the direction of the target UE’s uplink signal. Using these devices, the attacker aims to determine the physical location of the victim’s UE by measuring the direction of uplink signals at multiple points. These devices have the capability to passively monitor both uplink and downlink LTE signals. This capability can be achieved using an open-source LTE eavesdropper (e.g., LTESniffer (Hoang et al., 2023)) with software-defined radio devices (e.g., USRP (Research, 2014)). One unique capability of attackers is the ability to actively inject malicious LTE signals (i.e., signal overshadowing (Erni et al., 2022; Yang et al., 2019)) on both downlink and uplink channels. Upon obtaining the victim’s uplink scheduling information using these devices, the attacker can employ a directional antenna to measure the direction of the uplink signal source, detecting the strongest signal power. Utilizing AoA measurements from multiple locations, the attacker performs a multiangulation process to ascertain the victim’s precise coordinates.
Table 1. Uplink Tx power depending on the distance between the eNB and the UE.
Distance (m) 10 30 50 70 90 110
Tx PWR (dBm) -7.04 0.49 5.05 7.65 7.2 7.56
RSRP (dBm) -62.37 -73.43 -78.91 -85.36 -86.45 -87.6
  • RSRP: Reference Signals Received Power.
@@
Additionally, among various cellular localization techniques, we use AoA measurements of uplink signals as a fundamental technique. Other techniques such as TA, ToA, and TDoA have some limitations in employing fine-grained localization of uncooperative UEs in real-world environments. ToA-based cellular localization (Kotuliak et al., 2022) requires an additional process for fingerprinting the victim’s device to account for hardware errors. This is essential for accurately calculating the arrival signal timing, to precisely estimate the victim’s location. In addition, TA-based localization (Tan et al., 2021) shows only sub-cell-level localization performance with 78m granularity.
3.3. Challenges
@@
Physical localization of uncooperative devices in real-world environments needs to address the following challenges:
@@
C1: Dependency on victim’s uplink signal. Recall that the primary property exploited by localization attackers is monitoring uplink signals generated by the victim’s UE. From this perspective, the attacker can initiate localization only when the victim generates uplink traffic. In other words, localization is challenging when the victim is silent (i.e., not transmitting any signal). In such instances, the attacker is compelled to passively wait for the victim’s uplink transmission, as they lack access to either the eNB or victim UE.
@@
Additionally, it is important to note that RNTI is a temporary identifier that changes frequently (approximately every 15 to 30 seconds (Bae et al., 2022; Kohls et al., 2019; Rupprecht et al., 2019)). This implies that when a UE is not actively utilizing any services for a specific duration after connection, its radio connection is terminated by the eNB, and the previously assigned RNTI expires. If the UE later requests LTE services from the eNB, a new RNTI will be assigned for radio communication. The attacker may face a challenge in identifying the victim’s location if their radio connection and RNTI expire during the localization process. This necessitates the attacker to patiently wait for the victim to establish a new radio connection, after which they must track the victim’s new RNTI to resume the localization.
@@
C2: Undetectable victim’s signal. The successful measurement of arrival signals at the attacker’s sniffer is crucial for executing the localization attack. If the sniffer fails to detect the victim’s uplink signal, estimating the victim’s location becomes challenging. This issue arises because the UE does not consistently transmit its signal with high power to the eNB. Notably, when the UE is situated in close proximity to the eNB, weak uplink transmission power is used. The UE dynamically adjusts its uplink transmission power to compensate for signal path loss, influenced by factors such as distance and obstacles, in coordination with the eNB (3GPP. TS 36.213, v17.1.0, 2022; Michalevsky et al., 2015). For instance, Table 1 demonstrates how the UE adjusts its transmission power based on both distance and RSRP.
@@
The transmission power control between the eNB and UE creates a shadow area where the victim’s signal remains undetectable by the sniffer. To illustrate this shadow area in real-world environments, we examine how the UE’s position influences the sniffer’s signal detection in an operational commercial network. In our experiment, the sniffer, positioned 80m away from the target UE, successfully detected uplink signals from the UE located 90m away from the eNB. However, when the UE was only 10m away from the eNB, the sniffer failed to detect uplink signals at the same position (refer to our website (UMA, 2024) for detail). Unfortunately, due to the attacker’s lack of permission to control networks and UE, they are unable to legitimately instruct the UE to increase its transmission power when encountering the shadow area.
Figure 2. Impact of cellular repeater in localization.
Figure 3. Overview of UMA.
@@
C3: Differentiating victim from repeater. A cellular repeater is designed to amplify both downlink and uplink cellular signals, reducing signal path loss between the eNB and UE. As a consequence of the repeater’s operation, the UE tends to transmit uplink data with lower signal power. Concurrently, the repeater transmits (i.e., relays) the data with amplified signal power. It is widely deployed to enhance cellular service coverage. However, it poses a challenge for attackers in determining the direction of the strongest signal.
@@ 3.4. Approach: UMA
We present an Uncooperative Multiangulation Attack (UMA), which resolves the key challenges of tracking uncooperative devices (§ 3.3). UMA leverages AoA measurements with directional antennas, composed of six steps (Figure 3). These steps include RNTI acquisition using the victim’s online identity (①–②), manipulation of uplink scheduling (③), boosting the target’s uplink signal strength (④), and measurement of the AoA of the uplink signal (⑤–⑥). Note that while we assume authorities can obtain the RNTI with assistance from operators, as described in § 3.2, we include the RNTI acquisition process using online identity for the sake of completeness in explaining the overall tracking process in UMA.
@@
In particular, we come up with two approaches to effectively address the challenges of localizing uncooperative cellular devices; the scheduling manipulation attack and the power boosting attack. The scheduling manipulation attack is specifically designed to address challenge C1, allowing the attacker to acquire the unexpired victim’s RNTI and ensure continuous uplink transmission. To overcome challenges C2 and C3, we employ the power boosting attack, enabling the attacker to increase the victim’s uplink transmission power and distinguish the victim’s signal from the repeater’s. Both attacks exploit vulnerable protocols in LTE specification, which lack implemented security protection measures.
@@
Overall procedures of UMA. The first two steps are to obtain the victim’s RNTI associated with their phone number (§ 4.1). This is accomplished by sending silent Short Message Service (SMS) messages (①) and monitoring traffic patterns on the downlink channel (②). Next, the attacker proceeds with a scheduling manipulation attack (③); it transmits fake scheduling requests, and reports fake transmission buffer status to the eNB, posing as the victim UE with its RNTI (§ 4.2). This attack ensures the victim’s radio connection remains active, and their RNTI stays unexpired. Through the scheduling manipulation attack, the attacker forces the victim UE to generate uplink traffic even when there is no data to transmit in its current buffer. Subsequently, the attacker conducts a power boosting attack (④) by injecting manipulated commands for transmission power control logic into the victim’s UE, disguising as the eNB. This attack forces the victim UE to increase its uplink transmission power up to a maximum level (23dBm), overcoming practical challenges caused by the undetectable victim’s uplink signal and the presence of a cellular repeater in the radio network (§ 4.3). After these active operations, the attacker obtains the victim’s scheduling information broadcasted over DCI 0 by monitoring the downlink channel (⑤). With this information, the attacker identifies the victim’s signal on the uplink channel. Then, the attacker measures the AoA of the victim’s uplink signal by rotating a directional antenna mounted on the sniffer (⑥). Ultimately, the attacker can achieve physical localization by tracking down the victim’s location iteratively or employing multiangulation with multiple sniffers.
4. Design of UMA
@@
This section presents the design of our approaches. Firstly, we introduce a simple yet feasible method, inspired by previous works, for acquiring the victim’s RNTI using only their online identity (§ 4.1). Next, we describe how an unprivileged attacker compels the victim to generate uplink traffic and maintains the victim’s RNTI (§ 4.2). Finally, we illustrate how the attacker compels the victim to increase their uplink transmission power and overcomes cellular repeaters (§ 4.3).
4.1. From Online Identity to Radio Identifier
@@
For an attacker armed solely with the target’s online identity (e.g. phone number, WhatsApp ID), the initial step in UMA is to acquire the target’s temporary identifier used in the radio network (RNTI). The eNB broadcasts uplink scheduling information to each UE, designating a recipient UE by using the RNTI. Consequently, an unprivileged attacker seeking to monitor and ascertain the uplink signal of the victim UE must be aware of the victim’s RNTI. Since the RNTI is intentionally designed to be decoupled from the online identity, acquiring the RNTI becomes essential to establish the link between the victim’s online identity and its corresponding RNTI.
@@
Inspired by prior works (Hong et al., 2018; Hussain et al., 2019; Kohls et al., 2019; Kune et al., 2012; Ludant et al., 2023; Rupprecht et al., 2019; Shaik et al., 2016), our approach to acquiring the RNTI involves correlating voice/SMS traffic pattern directed toward the victim with monitored downlink packets identified by the recipient’s RNTI.
@@
RNTI acquisition is conducted through the following three steps. ➀ Given the victim’s online identity, the attacker successively generates silent voice calls or silent SMSes to the victim following a specific traffic pattern (e.g., multiple times with a constant time gap). ➁ The attacker monitors the downlink traffic of voice calls and SMSes. Specifically, they monitor the dedicated channels, called bearers, used for sending control messages of calls and SMSes. Note that the attacker can monitor the targeted traffic pattern clearly as each bearer is designed to deliver a different type of data. For example, messages for controlling voice calls are delivered over a bearer named Data Radio Bearer (DRB) 1, and user-plane traffic and SMS control messages are flooded via DRB 2 and Signalling Radio Bearer (SRB), respectively. ➂ The attacker examines radio connections and finds out the connection showing the intended traffic pattern. Finally, the attacker determines the RNTI of that connection as belonging to the victim.
4.2. Manipulating Uplink Scheduling
@@
One of the key pillars of UMA lies in its ability to a) compel the victim to generate uplink transmission and b) retain the acquired victim’s radio identifier (RNTI). To achieve this, we introduce a novel approach known as the Scheduling Manipulation Attack. Essentially, the Scheduling Manipulation Attack injects a crafted message into the eNB, prompting the eNB to schedule resources for uplink transmission to the victim UE and triggering the UE to transmit uplink data.
4.2.1. Uplink Scheduling Procedure
@@
We begin by outlining the control mechanism governing a UE’s uplink data transmission. Essentially, the UE is granted permission to transmit uplink data only when it receives uplink grants allocated by the eNB. The UE follows a defined procedure to secure uplink grants for data transmission to the eNB. Initially, the UE transmits a scheduling request (3GPP. TS 36.213, v17.1.0, 2022), signaling to the eNB its intention to transmit uplink data. Upon receiving the scheduling request, the eNB grants approval for the UE’s Buffer Status Reporting (BSR) (3GPP. TS 36.321, v16.6.0, 2021), a mechanism used to report the size of uplink data in its buffer. Subsequently, the UE initiates the BSR, leading to the allocation of sufficient uplink grants by the eNB, allowing the UE to transmit its data.
Figure 4. Procedure of scheduling manipulation attack.
@@
Scheduling Request is a PHY layer message sent from the UE to the eNB over the Physical Uplink Control Channel (PUCCH). To transmit a scheduling request to the eNB, the UE uses schedulingRequestConfig (3GPP. TS 36.213, v17.1.0, 2022; 3GPP. TS 36.331, v17.3.0, 2023), which was previously shared via the RRC connection setup message. This configuration notifies each UE which resource should be used to request scheduling.
@@
4.2.2. Design of Scheduling Manipulation Attack